### v1

# gather+[1][1]

Abstract.
Matricized Tensor Times Khatri-Rao Product (MTTKRP) is the computational bottleneck in sparse tensor decomposition. As real-world sparse tensors grow to billions of nonzeros, they increasingly demand higher memory capacity and compute throughput from hardware accelerators. In this work, we present AMPED, a multi-GPU parallel algorithm designed to accelerate MTTKRP on billion-scale sparse tensors. AMPED scales beyond the limits of a single GPU, meeting both the memory and performance requirements of large-scale workloads. We introduce a partitioning strategy combined with a dynamic load balancing scheme to distribute computation and minimize GPU idle time. On real-world billion-scale tensors, AMPED achieves a 5.1$\times$ geometric mean speedup in total execution time over state-of-the-art GPU baselines using 4 GPUs on a single CPU node.
MTTKRP, multi-GPU, Tensor Decomposition
copyright: nonejournalyear: 2025copyright: rightsretainedconference: 54th International Conference on Parallel Processing; September 8 - 11, 2025; San Diego, CAbooktitle: 54th International Conference on Parallel Processing (ICPP ’25), September 8 - 11, 2025, San Diego, CAccs: Computing methodologies Concurrent algorithmsccs: Computing methodologies Massively parallel algorithmsccs: Computing methodologies Vector / streaming algorithms
1. Introduction
Distribution Statement A: Approved for public release. Distribution is unlimited.
Tensors provide a natural way to represent data with multiple dimensions. Tensor decomposition transforms tensors with higher dimensionalities to a reduced latent space that can be leveraged to learn salient features of the underlying data distribution. Different application domains, including machine learning (Ji et al., 2019), signal processing (Cichocki et al., 2015), and network analysis (Fernandes et al., 2021) have employed tensor decomposition to achieve superior performance compared to conventional approaches. Canonical Polyadic Decomposition (CPD) (Hong et al., 2020) has become the widely used tensor decomposition method, where the Matricized Tensor Times Khatri-Rao Product (MTTKRP) is the computational bottleneck (Kolda and Bader, 2009).
Since real-world tensors are generally sparse, tensor formats only keep nonzero tensor elements to reduce memory consumption (Wijeratne et al., 2021; Helal et al., 2021; Wijeratne et al., 2023b). There is a need to develop optimized sparse tensor data layouts that support the highly irregular data access patterns of MTTKRP while accessing input tensors and factor matrices (Wijeratne et al., 2021; Helal et al., 2021; Wijeratne et al., 2023b).
Real-world tensors often exhibit irregular shapes and distributions of nonzero tensor elements, which pose significant challenges when performing MTTKRP computations on multiple GPUs. These challenges arise from irregular memory access patterns, load imbalance among many GPU streaming multiprocessors, and the synchronization overhead among GPUs.
The size of real-world tensors is increasing rapidly with recent advances in big data applications (Cichocki, 2014). Such applications use tensors with billions of nonzero tensor elements (i.e., billion-scale tensors). Therefore, parallel algorithms that scale beyond a single GPU are required to perform MTTKRP (Bharadwaj et al., 2024; Chen et al., 2022a) on such large tensors. However, distributing the sparse tensor across multiple GPUs can lead to load imbalance, latency in data migration, and intermediate value communication among devices, which can result in additional overhead to execution time.
Recent works have proposed accelerating MTTKRP on CPU (Helal et al., 2021; Kurt et al., 2022; Wijeratne et al., 2023a; Laukemann et al., 2024), GPU (Nisa et al., 2019a, c; Wijeratne et al., 2024; Nguyen et al., 2022a; Lin et al., 2024), and ASIC (Srivastava et al., 2020). In CPU and ASIC, the external memory is large enough to maintain billion-scale tensors. Meanwhile, general-purpose GPUs encounter additional scalability challenges on a single device due to limited GPU global memory.
In prior work, multiple GPUs were used to accelerate MTTKRP on millions-scale sparse tensors (Chen et al., 2022b). Such implementations do not scale well on billion-scale sparse tensors because of (1) the significant idle time incurred by some GPUs due to workload imbalance across GPUs, (2) increased communication overhead due to the latency in sharing intermediate results across GPUs, (3) the synchronization overhead across GPUs, and (4) scheduling overhead of dynamic load-balancing schemes during execution.
Various efforts have been made to balance the workload of MTTKRP across streaming multiprocessors (SMs) of a single GPU (Nisa et al., 2019c; Wijeratne et al., 2024; Nguyen et al., 2022a; Li et al., 2018a). FLYCOO-GPU (Wijeratne et al., 2024) introduces a partitioning scheme that eliminates task dependencies between GPU SMs for each mode and minimizes workload imbalance among GPU SMs. Our work extends the FLYCOO-GPU (Wijeratne et al., 2024) partitioning scheme to a multi-GPU environment, addressing a new set of challenges: (1) eliminating data dependencies between tensor partitions to minimize GPU-GPU communication and synchronization overhead, (2) balancing the workload across GPUs to minimize the idle time for each GPU, and (3) avoiding host CPU computations when collecting and distributing input tensor partitions and factor matrices, since CPU computing power is significantly lower than GPUs.
Table 1. Summary of Related Work
Work Number of tensor Multi-GPU Load-balancing Support for Task independent
copies required support billion-scale tensors partitioning across GPUs
AMPED (ours) No. of modes
MM-CSF (Nisa et al., 2019c) No. of modes
FLYCOO-GPU (Wijeratne et al., 2024) 2
BLCO (Nguyen et al., 2022a) 1
HPSPTM (Chen et al., 2022b) No. of modes
ParTI-GPU (Li et al., 2018a) 1
*: Only support single GPU
The key contributions of this work are:
  • We introduce a novel parallel algorithm to perform MTTKRP on sparse tensors using multiple GPUs. The proposed algorithm achieves geometric mean speedups of 1.9$\times$, 2.3$\times$, and 3.3$\times$ in execution time while using 2, 3, and 4 GPUs, compared to a single GPU implementation on billion-scale tensors.
  • We propose a tensor partitioning scheme to distribute the tensor elements across multiple GPUs. The proposed partitioning scheme achieves a geometric mean speedup of 8.2$\times$ in total execution time compared with equally distributing the nonzero tensor elements across the GPUs.
  • We introduce a load balancing scheme to distribute the computations of MTTKRP among GPUs while eliminating the communication of intermediate values across GPUs in each mode of computation. After each mode of computation, the updated factor matrix rows are shared across the GPUs. The proposed load balancing scheme shows less than $1\%$ overhead in computation time across GPUs on billion-scale tensors.
  • Our work achieves a geometric mean speedup of 5.1$\times$ in total execution time on billion-scale tensors compared with the state-of-the-art GPU baselines.
2. Background and Related Work
2.1. Introduction to Tensors
In this Section, we introduce terminology of tensors and tensor decomposition used in the paper. For a complete review of the tensors, see (Kolda and Bader, 2009).
A tensor is a multidimensional array whose dimensions are commonly referred to as modes. A $N$ mode tensor with real values is represented as $\mathcal{T}\in\mathbb{R}^{I_{0}\times\cdots I_{d}\times\cdots\times I_{N-1}}$, where the mode $d$ refers to the $d^{\text{th}}$ dimension of the tensor. Individual elements of the tensor, such as $x=\mathcal{T}(i_{0},i_{1},i_{2},...,i_{N-1})$, are indexed by $i_{0},i_{1},i_{2},...,i_{N-1}$, which are referred to as the indices of the nonzero tensor element. Tensors are often converted to a matrix form through a process known as matricization or matrix unfolding. The mode $d$ matricization of $\mathcal{T}$, denoted as $\mathcal{T}_{(d)}$, reorganizes the tensor into a matrix $\mathcal{T}_{(d)}\in\mathbb{R}^{I_{d}\times(I_{0}\cdots I_{d-1}I_{d+1}\cdots I_{N-1})}$. Here, the rows of the matricized tensor correspond to the indices of the mode $d$, while the columns are formed by iterating through all possible combinations of the indices in the remaining modes (Favier and de Almeida, 2014).
2.2. Canonical Polyadic Tensor Decomposition
In this Section, we illustrate CPD using a 3-mode tensor $\mathcal{T}\in\mathbb{R}^{I_{0}\times I_{1}\times I_{2}}$ as an example.
CPD aims to approximate a tensor as a sum of rank-1 tensors. Specifically, given a tensor $\mathcal{T}$, the goal is to express it as:
$\mathcal{T}\approx\sum_{r=0}^{R-1}\mathbf{f{0}}_{r}\circ\mathbf{f{1}}_{r}\circ\mathbf{f{2}}_{r},$
where $R$ is a positive integer, and $\mathbf{f{0}}_{r}\in\mathbb{R}^{I_{0}}$, $\mathbf{f{1}}_{r}\in\mathbb{R}^{I_{1}}$, and $\mathbf{f{2}}_{r}\in\mathbb{R}^{I_{2}}$ for $r=0,\ldots,R-1$. Vectors $\mathbf{f{0}}_{r}$, $\mathbf{f{1}}_{r}$, and $\mathbf{f{2}}_{r}$ can be organized as factor matrices, $\mathbf{F_{0}}=[\mathbf{f{0}}_{0},\ldots,\mathbf{f{0}}_{R-1}]$, $\mathbf{F_{1}}$, and $\mathbf{F_{2}}$, respectively.
The rank of a tensor $\mathcal{T}$ is defined as the number of rank-1 tensors needed to exactly represent $\mathcal{T}$. When the rank is $R$, the CPD with $R$ components is computed by minimizing the Frobenius norm of the difference between the tensor and its approximation:
$\min_{\hat{\mathcal{T}}}\quad$ $\lVert\mathcal{T}-\hat{\mathcal{T}}\rVert^{2}_{F},$
subject to $\hat{\mathcal{T}}=[\![\lambda;\mathbf{F_{0}},\mathbf{F_{1}},\mathbf{F_{2}}]\!]=\sum_{r=0}^{R-1}\lambda_{r}\cdot\mathbf{f{0}}_{r}\circ\mathbf{f{1}}_{r}\circ\mathbf{f{2}}_{r},$
where $[\![\lambda;\mathbf{F_{0}},\mathbf{F_{1}},\mathbf{F_{2}}]\!]$ represents the Kruskal form of the tensor, and $\lambda_{r}$ is the weight of the $r$-th component.
The Alternating Least Squares (ALS) algorithm is the most commonly used method to compute the CPD due to its simplicity and computational efficiency. A critical kernel in CP-ALS is the Matricized Tensor-Times Khatri-Rao Product (MTTKRP), which dominates the computational cost. CP-ALS updates the factor matrices during each iteration by computing MTTKRP for each mode while keeping the other factor matrices fixed. For a 3-mode tensor, the MTTKRP operations are expressed as:
$\mathbf{\mathbf{F_{0}}}=\mathcal{T}_{(0)}(\mathbf{F_{1}}\odot\mathbf{F_{2}}),$
$\mathbf{\mathbf{F_{1}}}=\mathcal{T}_{(1)}(\mathbf{F_{2}}\odot\mathbf{F_{0}}),$
$\mathbf{\mathbf{F_{2}}}=\mathcal{T}_{(2)}(\mathbf{F_{0}}\odot\mathbf{F_{2}}),$
In this work, we focus on MTTKRP on sparse tensors, where the tensor $\mathcal{T}$ is sparse and the factor matrices $\mathbf{F_{0}}$, $\mathbf{F_{1}}$, and $\mathbf{F_{2}}$ are dense. Sparse MTTKRP is particularly challenging due to the irregular structure of the tensor, which demands efficient handling of sparsity to optimize performance.
2.3. Related Work
FLYCOO-GPU (Wijeratne et al., 2024) proposes a single GPU-based parallel algorithm to accelerate MTTKRP on sparse tensors while adopting the FLYCOO tensor format (Wijeratne et al., 2023b). FLYCOO-GPU (Wijeratne et al., 2023b) introduces dynamic tensor remapping on the GPU to reorder the tensor during execution time, allowing mode-specific optimizations. Dynamically remapping the input tensor limits the scalability of FLYCOO-GPU across GPUs as the inter-GPU communication becomes the bottleneck in multiple GPUs. Unlike (Wijeratne et al., 2024), our work supports multiple GPUs while avoiding inter-GPU task dependencies that lead to race conditions across GPUs.
BLCO (Nguyen et al., 2022a) proposes the blocked linearized coordinate format (BLCO) that enables out-of-memory computation where the input tensor is stored in host CPU external memory and streamed to a single GPU during the execution time of each mode computation. The streaming data-based computations of BLCO enable the partitions of the input tensor to be stored inside a large host CPU external memory, which is loaded into a single GPU over time and performs the MTTKRP partition by partition. Unlike BLCO, our approach distributes the computations across several GPUs while balancing the workload among all the GPUs.
HPSPTM (Chen et al., 2022b) proposes a framework to exploit multilevel parallelism and data reusability in heterogeneous HPC systems, including CPU + multi-GPU platforms. The data scheduling and partitioning scheme of HPSPTM dynamically transfers tensor partitions, intermediate results, and rows of factor matrices between GPUs during each factor matrix computation. In contrast, our work introduces a static load balancing scheme to minimize the data transfers across GPUs during each factor matrix computation. HPSPTM focuses on million-scale tensors, while our work focuses on billion-scale tensors.
MM-CSF (Nisa et al., 2019a) and HiCOO-GPU (Li et al., 2018a) propose novel tensor formats to distribute workload on a single GPU. In contrast, we propose a parallel algorithm to perform MTTKRP using multiple GPUs, which requires optimizations to reduce communication and synchronization overheads across GPUs.
Table 1 compares the characteristics of the related work with our work. Unlike related work, our work supports billion-scale sparse tensors. Our work also uses a task-independent partitioning scheme that balances the workload among the GPUs and their streaming multiprocessors.
3. Tensor Partitioning Scheme
In this work, we extend the partitioning scheme proposed in FLYCOO-GPU (Wijeratne et al., 2024) to distribute the nonzero tensor elements among multiple GPUs. Specifically, we adopt the FLYCOO-GPU partitioning scheme to distribute the tensor across GPUs, leveraging its task independence property, which allows each tensor partition to operate independently of the other partitions. For completeness, we describe the adopted partitioning scheme as inter-device partitioning.
Unlike FLYCOO-GPU (Wijeratne et al., 2024), our approach does not use dynamic tensor remapping. Consequently, we avoid embedding shard IDs within each nonzero tensor element. Instead, we maintain multiple copies of the input tensor in CPU external memory. For a detailed explanation of dynamic tensor remapping and the use of shard IDs, please refer to (Wijeratne et al., 2024).
Furthermore, we use the same notation introduced by FLYCOO-GPU (Wijeratne et al., 2024): When performing MTTKRP for the mode $d$ of an input tensor, we denote the mode $d$ as the output mode and its corresponding factor matrix as the output factor matrix. The rest of the tensor modes are called input modes, and the corresponding factor matrices are called input factor matrices.
3.1. Output Factor Matrix Row Update Computation (OFMRU)
Updating an output factor matrix row is the fundamental operation of MTTKRP. During MTTKRP along each mode (see Section 2.2), the output factor matrix rows are updated using each nonzero tensor element of the input tensor.
[…]
{gather+}
[0.85]  F_d(i_d, r) = F_d(i_d, r) + val(X(i_0, …, i_d, …, i_N-1)) ⋅∏_k = 0 k ≠d^N-1 F_k(i_k, r)
Following Section 2, Equation 3.1 shows the computation to update the output factor matrix, $\mathbf{F_{d}}$ using nonzero tensor element, $\mathcal{X}(i_{0},..,i_{d},..,i_{N-1})$ of a $N$ mode tensor. Here, $i_{0},..,i_{d},..,i_{N-1}$ is the position of the nonzero tensor element and $val(\mathcal{X}(i_{0},..,i_{d},..,i_{N-1}))$ is the value of the nonzero tensor element. Equation 3.1 computes the value in each column $r$ ($0\leq r<R$) in row $i_{d}$ of the mode $d$ factor matrix ($0\leq d<N$). Note that the row update is performed on all the nonzero tensor elements of the input tensor $\mathcal{X}$ to compute MTTKRP of output mode $d$.
3.2. Tensor Partitioning
We introduce a static partitioning scheme to partition an input tensor along each mode of an input tensor. Similarly to related works (Nisa et al., 2019c, a), we use multiple copies of the input tensor, where each tensor copy is partitioned targeting an output mode computation. Note that all tensor copies are stored in the large external memory of the host CPU. Our partitioning scheme in each mode distributes the nonzero tensor elements across (1) multiple GPUs using inter-device partitioning and (2) streaming multiprocessors (SMs) within each GPU using inter-SM partitioning. Note that an inter-device partition is executed by a single GPU, and an inter-SM partition is executed by a single GPU SM.
3.2.1. Inter-device Partitioning
The proposed partitioning scheme organizes the nonzero tensor elements of the input tensor into inter-device partitions (iDp) based on the output mode index along each mode. For a given output mode $d$, the OFMRU introduced in Section 3.1 (see Equation 3.1) is performed on each nonzero tensor element to update the corresponding rows of the output factor matrix. All nonzero tensor elements sharing the same output factor matrix index contribute to updating the same row. GPUs should maintain coherence among the updates that use nonzero tensor elements with the same output mode indices to avoid race conditions. To prevent race conditions between GPUs, all nonzero tensor elements sharing the same output mode index are assigned to the same $iDp$, eliminating task dependencies across GPUs. It avoids the need to maintain coherency across GPUs during execution, thereby significantly reducing synchronization overhead and inter-GPU data communication.
3.2.2. Inter-SM Partitioning
The nonzero tensor elements of each iDp are equally distributed among the SMs of each GPU by partitioning each iDp into equal sized inter-SM partitions (iSMp). With iSMp partitions, all the SMs of a GPU are assigned the same workload during execution. We use atomic operations to avoid race conditions between SMs of the same GPU.
3.3. Tensor Format Definition
Consider a multi-GPU setup with $m$ GPUs, each containing $g$ GPU SMs. For each output mode $d$, we divide the output mode indices $I_{d}$ into sets of equal sized partitions $I_{d,0},I_{d,1},\ldots,I_{d,k_{d}-1}$, where $k_{d}=\frac{|I_{d}|}{m}$. Here, $|I_{d}|$ denotes the size of $I_{d}$. Each index partition $I_{d,j}$ ($j=0,1,\ldots,(k_{d}-1)$) is a subset of the output mode indices $I_{d}$. Next, all nonzero tensor elements that incident on the indices in $I_{d,j}$ are collected into an inter-device partition, denoted by $\textit{{iDp}}_{d,j}$. To distribute nonzero tensor elements among GPU SMs, we further divide each inter-device partition into equal-sized sets called inter-SM partitions (iSMp). Each inter-device partition, $\textit{{iDp}}_{d,j}$ is divided into $t_{d,j}=\left\lceil|\textit{{iDp}}_{d,j}|/g\right\rceil$ inter-SM partitions. We denote the $q$-th inter-SM partition in $\textit{{iDp}}_{d,j}$ as $\textit{iSMp}_{d,j,q}$. For a tensor with $|\mathcal{T}|$ nonzero tensor elements, the total number of inter-SM partitions in mode $d$ is $\tau_{d}=\sum_{h=0}^{k_{d}-1}t_{d,h}\approx\frac{|\mathcal{T}|}{(m\times g)}$.
Observation 1.
Inter-device partitions (iDps) of a given output mode can be executed independently of other inter-device partitions of the same mode.
Proof.
Consider a nonzero tensor element $x_{i}\in\textit{iDp}_{d,j}$, where $\textit{iDp}_{d,j}$ is the $j^{\textrm{th}}$ inter-device partition of the input tensor $\mathcal{T}$ in mode $d$. Let the output mode index of $x_{i}$ in mode $d$ be $c_{d}$, meaning that $x_{i}$ contributes to updating the $c_{d}^{\textrm{th}}$ row of the output factor matrix of mode $d$ during the output mode $d$ MTTKRP computation.
From Equation 3.1 in Section 3.1, $x_{i}$ has data dependencies with any other tensor element $x_{z}$ if $x_{z}$ also updates the $c_{d}^{\textrm{th}}$ row of the output factor matrix. This implies that $x_{i}$ and $x_{z}$ share the same output mode index, $c_{d}$, and therefore contribute to the same output factor matrix row.
According to the tensor partitioning scheme defined in Section 3.2, all nonzero tensor elements with the same output mode index $c_{d}$ are assigned to the same iDp partition, $\textit{iDp}_{d,j}$. Thus, any $x_{z}$ with the same output mode index $c_{d}$ must also belong to $\textit{iDp}_{d,j}$ ($x_{z}\in\textit{iDp}_{d,j}$). Consequently, all nonzero tensor elements that update the same row of the output factor matrix are guaranteed to be in the same partition.
Since no two partitions share nonzero tensor elements with the same output mode index, there are no data dependencies between different iDp partitions of mode $d$. As a result, all $\textit{iDp}_{d,j}$ for a given mode $d$ can be executed independently. ∎
With Observation 1, iDps of each output mode can be executed on any GPU independently of each other.
4. Parallel Algorithm
4.1. CUDA Programming Model
In the CUDA programming model (Owens et al., 2008; Hijma et al., 2023), a multi-threaded program is partitioned into blocks of threads (i.e., threadblocks) where each threadblock operates independently. The threadblocks are organized into a GPU Grid (Ansorge, 2022; Zeller, 2011). A multi-GPU implementation executes a kernel as Grids of threadblocks where each Grid is executed on a separate GPU. Each threadblock is executed by one GPU streaming multiprocessor (SM).
4.2. Tensor Partition and Threadblock Mapping
Following the partition scheme proposed in Section 3.2, the inter-device partitions (iDp) and the inter-SM partitions (iSMp) are assigned to GPU Grids and threadblocks, respectively (Ansorge, 2022), as illustrated in Figure 1. Figure 1 uses the same notation introduced in Section 3.2. Once a GPU finishes executing all the computations in a Grid, a new Grid is loaded to the GPU for execution. Similarly, when a GPU SM finishes executing all the computations in a threadblock, a new threadblock in the same Grid is assigned to the SM.
Figure 1. Tensor partition mapping of mode d
4.3. Target Platform
Figure 2. Target Platform
We consider a single node multi-GPU platform where the GPUs are connected to a single CPU, as illustrated in Figure 2. All the GPUs are connected using GPU to GPU interconnection (i.e., GPUDirect P2P (NVIDIA Corporation, 2024a)). Using the Cuda programming model, our parallel algorithm directly transfers data from one GPU global memory to another GPU global memory when required (see Section 4).
In Figure 2, we show GPU to GPU interconnection and host CPU to GPU interconnection separately for clarity. The host CPU to GPU interconnection and GPU to GPU interconnection can share the same physical hardware connections such as PCIe.
4.4. Data Distribution Among GPUs
All the tensor partitions (of all output modes) are stored in the host CPU external memory. When an $iDp$ partition of the tensor is ready for execution, $iDp$ is transferred to the global memory of the corresponding GPU.
The factor matrices of each billion-scale tensor are significantly smaller (i.e., few mega-bytes) than the GPU global memory. Each GPU maintains a local copy of the factor matrices in its global memory. Once the MTTKRP computation for an output mode is completed, the updated rows of the output factor matrix are exchanged between GPUs to prepare for the computation of the following output mode.
4.5. Overall Algorithm
1 Input: Input tensor copies, $\textbf{T}=\{\mathcal{T}_{(0)},\mathcal{T}_{(1)}\cdots\mathcal{T}_{(N-1)}\}$
2 Factor matrices from previous iteration, $\textbf{F}=\{F_{0},F_{1},...,F_{N-1}\}$
3 Output: Updated factor matrices $\hat{\textbf{F}}$
4 for each mode $d=0,1,\ldots,(N-1)$ do
5       // $\mathcal{T}_{in}$ takes the partitions corresponding to $d^{\text{th}}$ matricization of the input tensor
6       $\mathcal{T}_{in}\leftarrow\mathcal{T}_{(d)}$
7       // Execute Grids using multiple GPUs
8       for each inter-device partition, $iDp_{d,z}$ in $\mathcal{T}_{in}$, $z=0,1\ldots,k_{d}-1$ parallel do
9             // Execute threadblocks in parallel across GPU SMs
10             for each inter-SM partition $iSMp_{d,z,b}$ in $iDp_{d,z}$, $b=0,1,\ldots,t_{d,z-1}$ parallel do
11                   $nnz\leftarrow 0$
12                   for $nnz<|iSMp_{d,z,b}|$ parallel do
13                         for each column, $l$ in threadblock parallel do
14                               if $nnz+l<|iSMp_{d,z,b}|$ then
15                                     Load($x_{i}$ at $(nnz+l)$)
16                                     $\hat{F}_{d,z}\leftarrow$OFMRU($x_{i}$,F)
17
18
19                        // $\theta$ = number of columns in a threadblock
20                         $nnz\leftarrow nnz+\theta$
21
22
23            __Inter-GPU Barrier__
24             // Update output factor matrix copy on each GPUs
25             // See Algorithm 3 on GPU to GPU communication
26             $F_{d}\leftarrow\hat{F}_{d,z}$
27
28      __Inter-GPU Barrier__
29
Algorithm 1 Overall proposed algorithm on multi-GPUs for a single iteration
In this Section, we present the parallel algorithm for a single iteration of tensor decomposition. In tensor decomposition, the proposed algorithm is iteratively performed to generate the factor matrices that best approximate the original input tensor.
Algorithm 1 shows the overall parallel algorithm to execute MTTKRP across all modes of a single iteration. Algorithm 1 takes (1) all copies of the $N$ mode input tensor $\textbf{T}=\{\mathcal{T}_{(0)},\mathcal{T}_{(1)}\cdots\mathcal{T}_{(N-1)}\}$ and (2) the factor matrices $\textbf{F}=\{F_{0},F_{1},...,F_{N-1}\}$ computed in the previous iteration. MTTKRP is performed mode by mode (Algorithm 1: line 4). Each GPU operates on the inter-device partitions ($iDps$) statically assigned by the proposed load balancing scheme in Section 4.9 prior to computation.
Once an $iDp$ is assigned to a GPU for computation, the GPU operates on $iDp$ independently of other GPUs (Algorithm 1: line 8). Note that $iDps$ do not have task dependencies between them and can be executed independently (see Observation 1). Each GPU SM of the GPU operates on an inter-SM partition ($iSMp$) in the assigned $iDp$ (Algorithm 1: line 9-10).
In Algorithm 1, $iSMp_{d,z,b}$ denotes the $b^{\text{th}}$ inter-SM tensor partition of the $z^{\text{th}}$ $iDp$ in mode $d$ (see Section 3.2). When a GPU SM is available for computation, a threadblock and its associated inter-SM tensor partition are assigned to the GPU SM. The threadblock performs OFMRU on each nonzero tensor element in the assigned $iSMp$ (Algorithm 1: line 18). To support OFMRU, we use a threadblock with dimensions of $R\times\theta$ executed on each GPU SM. Here, $R$ denotes the rank of factor matrices, and $\theta$ indicates the number of nonzero tensor elements loaded to a threadblock at once. $R\times\theta$ is kept equal to the number of threads in a GPU SM to utilize all the threads in a GPU SM. Each column in the threadblock loads a nonzero tensor element at a time and distributes it across the threads in the same column. Each thread in the column extracts the embedded information from the loaded nonzero tensor element (Algorithm 1: lines 12-18). and performs OFMRU (Algorithm 1: line 16).
After processing $iDp$, the updated factor matrix rows are shared across the GPUs (Algorithm 1: line 22) and the local copy of each GPU is updated before executing MTTKRP for the next output mode.
1 OFMRU($x_{i}$, F):
2 Input: Nonzero tensor element $x_{i}$
3           Factor matrices $\textbf{F}=\{F_{0},F_{1},...,F_{N-1}\}$
4 Output: Updated $F_{d,z}$
5 $value\leftarrow val_{i}$
6 $\beta_{i}=(c_{0},\ldots,c_{N-1})$
7
8// $\ell$ is a vector of size R
9 for each rank $r=0,\ldots,R-1$ parallel do
10       $\ell(r)\leftarrow value$
11
12for input mode $w\in\{0,\ldots,N-1\}\setminus\{d\}$ do
13
14      $vec\leftarrow$ Load(row $c_{w}$ from $w^{\text{th}}$ factor matrix, $F_{w}$)
15       // Row 00 to $R-1$ of the threadblock perform independent computations
16       for each rank $r=0,\ldots,R-1$ parallel do
17             $\ell(r)\leftarrow\ell(r)\times vec(r)$
18
19
20
21for each rank $r=0,\ldots,R-1$ parallel do
22       $Y_{d}(c_{d},r)\leftarrow\text{Atomic}(Y_{d}(c_{d},r)+\ell(r))$
23
Algorithm 2 Update the factor matrix row of mode $d$
4.6. Parallel Algorithm of Output Factor Matrix Row Update Computation
Algorithm 2 describes the Output Factor Matrix Row Update Computation (OFMRU) on each nonzero tensor element. In Algorithm 2, the rows of the input factor matrices are loaded from the global memory of the GPU (Algorithm 2: line 11) depending on the indices of the current nonzero tensor element ($x_{i}$) in the GPU thread. Subsequently, the SM threads perform OFMRU (see Section 3.1) using the nonzero tensor element and the rows of the factor matrices (Algorithm 2: lines 10 - 16). To enable threadwise parallelism in OFMRU, each thread in a column performs computations on its respective rank (Algorithm 2: lines 12 - 14). Each GPU threadblock updates the output factor matrix (Algorithm 2: lines 16), using atomic computations to avoid race conditions between threadblocks in the same GPU.
4.7. Host CPU-GPU Communication
The input tensor copies are initially stored in the host CPU memory. The inter-device partitions (iDps) are transferred from the host CPU memory to each GPU global memory during the execution time of each output mode.
4.8. GPU to GPU Communication
1 P2P Communication ($gpu\_id$, $F_{d,gpu_{i}d}$):
2 // Using ring network communication model
3 for each GPU $z=0,\ldots,(M-2)$ do
4       // Send and Receive commands are executed in parallel
5       Send_Copy = $F_{d,(gpu\_id+z)modM}$
6       Send(Send_Copy, $(gpu\_id+1)modM$, size(Send_Copy))
7       Rec_Copy = $F_{d,(gpu\_id-z-1)modM}$
8       Receive(Rec_Copy, $(gpu\_id-1)modM$, size(Rec_Copy)))
9       Barrier
10
Algorithm 3 GPU to GPU peer-to-peer communication
All GPUs maintain a local copy of the factor matrices in GPU global memory (see Section 4.4). As mentioned in Section 4.4, the factor matrices of each billion-scale tensor are significantly smaller (i.e., a few megabytes) than the GPU global memory. Hence, keeping a local copy of the factor matrices does not consume significant additional memory on each GPU.
For a given output mode, once a GPU completes processing the iDp assigned to it, the GPU communicates the updated rows to all the GPUs in the platform. The distributed factor matrix will act as an input factor matrix in the following modes of MTTKRP computations.
We adopt the ring network communication model (Li et al., 2018b) to broadcast the partitions of the factor matrices among the GPUs, as shown in Algorithm 3. We use GPU to GPU communication to send and receive partitions of the factor matrices across GPU without involving the CPU host memory. The ring network model is suitable for bulk transfers among neighboring devices with limited bandwidth, which is ideal for communicating the output factor matrix among multiple GPUs.
4.9. Load Balancing
After assigning an iDp to a GPU, the number of OFMRU operations performed by the GPU on the iDp is proportional to the number of nonzero tensor elements in the iDp. The number of nonzero tensor elements in each iDp varies according to the sparsity of the tensor. In our work, the workload is distributed by assigning iDps of different sizes to GPUs, aiming to minimize the difference between the number of nonzero elements assigned to each GPU for computation.
Consider an input tensor divided into iDps $\{\textit{iDp}_{d,j}:\forall\,j\}$ for an output mode $d$. Let $\eta_{d,j}$ denote the number of nonzero tensor elements in $\textit{iDp}_{d,j}$. We arrange iDps so that $\eta_{d,j}\geq\eta_{d,j^{\prime}}$ if $j<j^{\prime}$, thus ordering the iDps in descending order according to the number of nonzero tensor elements. Each iDp is iteratively assigned to the GPU with the fewest number of nonzero tensor elements assigned at the time. This procedure is repeated for all modes of the tensor.
For a given output mode, let $\eta_{\textrm{max}}$ be the largest number of nonzero tensor elements assigned to a single GPU and $\eta^{*}_{\textrm{max}}$ be the maximum number of nonzero tensor elements assigned to a GPU in the optimal distribution possible for the given iDps. Then our proposed solution guarantees $\eta_{\textrm{max}}\leq 4/3\cdot\eta^{*}_{\textrm{max}}$ (Wijeratne et al., 2023b; Graham, 1969).
5. Experimental Results
5.1. Experimental Setup
5.1.1. Platform
Figure 3. Total execution time (speedup of our work compared to each baseline is shown at the top of each bar)
We conducted our experiments on a single CPU node with multi-GPUs. The CPU node has 4 NVIDIA RTX 6000 Ada Generation GPUs and AMD EPYC 9654 host CPU. The GPUs are connected to the host CPU via PCIe. GPUs use GPUDirect Peer-to-Peer (P2P) (NVIDIA Corporation, 2024a) to communicate with each other. Note that NVIDIA RTX 6000 Ada GPUs do not support NVLink (NVIDIA Corporation, 2024b) for direct GPU communication. Hence, we use the Direct P2P communication in our experiments.
GPU Specification: NVIDIA RTX 6000 Ada Generation GPU features the Ada Lovelace GPU architecture with 142 Streaming Multiprocessors (SMs) and 18176 cores, sharing 48 GB of GDDR6 global memory. Note that the NVIDIA RTX 6000 Ada Generation GPU has more computing power and global memory than the NVIDIA A100 GPU.
Host CPU Specification: In our experiments, we use a 2-socket AMD EPYC 9654 CPU as the host CPU platform. Each AMD EPYC 9654 consists of 96 physical cores (192 threads) running at a frequency of 2.4 GHz, sharing 1.5 TB of CPU external memory.
Each GPU is connected to the host CPU through a PCIe interface with 64 GB/s data bandwidth.
5.1.2. Implementation
The source code is developed using CUDA C++ (Zeller, 2011) and compiled using CUDA version 12.2 (Fatica, 2008).
5.1.3. Datasets
We used all publicly available billion-scale tensors from the Formidable Repository of Open Sparse Tensors and Tools (FROSTT) dataset (Smith et al., 2017) and Recommender Systems and Personalization Datasets (Rappaz et al., 2021; McAuley, 2021). Table 2 summarizes the characteristics of the tensors.
Table 2. Characteristics of the sparse tensors
Tensor Shape Number of
Tensor Elements
Amazon (Smith et al., 2017) $4.8M\times 1.8M\times 1.8M$ $1.7B$
Patents (Smith et al., 2017) $46\times 239.2K\times 239.2K$ $3.6B$
Reddit-2015 (Smith et al., 2017) $8.2M\times 177K\times 8.1M$ $4.7B$
Twitch (Rappaz et al., 2021; McAuley, 2021) $15.5M\times 6.2M\times 783.9K\times 6.1K\times 6.1K$ $0.5B$
5.1.4. Baselines
We evaluate the performance of our work by comparing it with the state-of-the-art GPU implementations, BLCO (Nguyen et al., 2022a) (out-of-memory computation enabled), MM-CSF (Nisa et al., 2019c), HiCOO-GPU (Li et al., 2018a), and FLYCOO-GPU (Wijeratne et al., 2024). The out-of-memory computation of BLCO stores the partitions of the input tensor inside the host CPU external memory, which is loaded into a single GPU over time and performs the MTTKRP partition by partition. To achieve optimal results with HiCOO-GPU, we use the recommended configurations provided in the source code (Li et al., 2019). For our experiments, we utilize the open-source BLCO repository (Nguyen et al., 2022b), ParTI repository (Li et al., 2019), and MM-CSF (Nisa et al., 2019b) repository.
5.1.5. Default Configuration
In our experiments, we used 4 GPUs as the default number of GPUs in the system. We set $\theta=32$ and $R=32$ as the configuration of the threadblocks described in Section 4.5.
5.1.6. Performance Metric - Total Execution Time
Similar to the literature (Nisa et al., 2019c; Wijeratne et al., 2024; Nguyen et al., 2022a; Li et al., 2018a), we measure the performance using the execution time to compute MTTKRP across all modes of an input tensor in a single iteration of tensor decomposition.
5.2. Overall Performance
Figure 3 shows the total execution time of our work on 4 NVIDIA RTX 6000 Ada GPUs. The speedup achieved by our work compared with each baseline in each input tensor is displayed at the top of the respective bar. The runtime error indicates that the host CPU operating system terminated the baseline during execution due to insufficient memory on the target hardware to store the input tensor, factor matrices and intermediate values.
For evaluation, we set the rank of the factor matrices ($R$) to 32, similar to the state-of-the-art (Nisa et al., 2019c; Nguyen et al., 2022a; Li et al., 2018a). Our work demonstrates a geometric mean speedup of 5.1$\times$ compared with the state-of-the-art baselines.
When out-of-memory computation is enabled, BLCO (Nguyen et al., 2022a) stores the input tensor in the host CPU memory and loads the tensor partition by partition from the external memory of the host CPU to the GPU global memory. In BLCO, using a single GPU introduces additional memory traffic between the host CPU and the GPU. Meanwhile, our work has more effective bandwidth between the host CPU and the GPU since multiple GPUs can concurrently communicate with the host CPU. Our work shows a geometric speedup of 5.1×5.1$\times$5.1 × using 4 GPUs compared to BLCO.
MM-CSF (Nisa et al., 2019c) performs MTTKRP only on the Amazon dataset. For Patents and Reddit, the GPU ran out of memory during the execution time. ParTI-GPU (Li et al., 2018a) can perform MTTKRP on Amazon and Patents. Also, MM-CSF (Nisa et al., 2019c) and ParTI-GPU (Li et al., 2018a) do not support Twitch, which has 5 modes.
FLYCOO-GPU (Wijeratne et al., 2024) does not support Amazon, Patents, and Reddit since these tensors do not fit in the GPU global memory. FLYCOO-GPU requires maintaining 2 copies of the tensor in the GPU global memory. On Twitch, FLYCOO-GPU outperforms our work by 3.9$\times$ due to the communication overhead of our work. Twitch is the smallest billion-scale tensor in the literature. The small size enables keeping 2 tensor copies inside the global memory of a single GPU. Since FLYCOO-GPU only targets a single GPU, FLYCOO-GPU does not require GPU-GPU communication or host CPU-GPU communication.
Figure 4. Sustained compute performance on different number of GPUs
5.3. Sustained Compute Performance
Figure 4 shows the sustained compute performance of our work on different numbers of GPUs. In all tensors, we observe a nearly linear increase in sustained compute performance since the computation time (i.e., each GPU SM performing OFMRU and GPU global memory accesses) is the dominating factor in execution (see Section 5.5).
5.4. Impact of Load Balancing Scheme
Figure 5. Impact of proposed load balancing scheme
Section 4.9 describes the load balancing scheme used in this work to distribute the inter-device partitions among GPUs.
An alternative approach is to distribute the non-zero tensor elements equally among all GPUs. It introduces additional computations on the host CPU to merge the partial results of each inter-device partition. Figure 5 compares our proposed strategy with the equal distribution of nonzero tensor elements among GPUs. Our proposed load balancing scheme achieves 5.3$\times$ to 10.3$\times$ speedups in total execution time.
5.5. Execution Time Breakdown
Figure 6. Execution time breakdown of the input tensors
For large tensors such as Patents and Reddit, moving inter-device partitions from CPU host memory to GPU global memory is the major contributor to the total communication time. Tensors with a large number of indices (e.g., Amazon and Twitch) require frequent GPU-GPU communication to update factor matrices at the end of each output mode computation, which significantly contributes to the total communication time. In particular, Reddit exhibits a significant communication overhead ($32\%$) during the execution due to (1) the large number of nonzero tensor elements in the tensor and (2) the modes with a large number of indices.
5.6. Load Balancing Workload among GPUs
Figure 7. Computation time overhead among GPUs
As discussed in Section 4, each GPU updates a set of rows of the output factor matrix. In this section, the computation time of a GPU is defined as the total time to perform OFMRU on all nonzero tensor elements assigned to a GPU across all modes. The computation time overhead is the difference between the maximum and minimum computation times among GPUs on the target platform.
Figure 7 shows the computation time overhead among GPUs (as a percentage) on our target 4 GPU platform. To determine the computation time of each GPU, we execute each GPU grid (see Section 4.2) separately and measure the computation time in each output mode. The computation time overhead for all billion-scale tensors is less than $1\%$ (as a percentage of the total time required to perform all OFMRU computations using all 4 GPUs).
However, Twitch has the most computation time overhead due to some indices of the tensor corresponding to popular streamers and games in the Twitch platform, which leads to a disproportionately large number of nonzero tensor elements assigned to those indices. It results in a workload imbalance between GPUs, leading to a larger overhead.
5.7. Scalability
Figure 8 shows the speedup of each input tensor as the number of GPUs increases from 1 to 4. The speedup is calculated by dividing the total execution time of the input tensor on a single GPU by its total execution time in each case. Our proposed parallel algorithm achieves geometric mean speedups of 1.9$\times$, 2.3$\times$, and 3.3$\times$ when using 2, 3, and 4 GPUs, respectively, compared to a total execution time of a single GPU. The speedup increases nearly linearly with the number of GPUs for each tensor.
In the single GPU implementation, tensor partitions are stored in the host CPU external memory and loaded onto the GPU one iDp at a time in each mode.
Figure 8. Scalability of the proposed Algorithm
5.8. Preprocessing Time
The preprocessing of an input tensor involves generating tensor partitions following Section 3. Note that our work does not focus on the acceleration of preprocessing time. We have included a comparison of preprocessing times in Figure 9 for completeness. For comparison, we used the preprocessing time of BLCO. The host CPU described in Section 5.1.1 is used for preprocessing.
Figure 9. Preprocessing time
6. Conclusion and Future Work
In this paper, we proposed AMPED, a novel parallel algorithm designed to accelerate MTTKRP on billion-scale tensors using multiple GPUs. Our approach introduced a partitioning scheme that allowed each tensor partition to be executed independently of other partitions in each output mode. The partitioning scheme, coupled with the proposed load-balancing strategy to distribute the workload across all the GPUs, minimized the GPU idle time. AMPED achieved a geometric mean speedup of 5.1$\times$ (using 4 GPUs) in total execution time compared with the state-of-the-art GPU baselines.
In our future work, we will adapt the proposed parallel algorithm to heterogeneous computing platforms with different devices, such as multiple CPUs, GPUs, and Field Programmable Gate Arrays (FPGAs), to show the adaptability of the proposed parallel algorithm.
Acknowledgements.
This work is supported by the National Science Foundation (NSF) under grant OAC-2209563, CSSI-2311870, and in part by the DEVCOM Army Research Lab under grant W911NF2220159.