### v2

# AMPED: Accelerating MTTKRP for Billion-Scale Sparse Tensor Decomposition on Multiple GPUs

Abstract.
[…]
MTTKRP, multi-GPU, Tensor Decomposition
journalyear: 2025copyright: usgovmixedconference: 54th International Conference on Parallel Processing; September 08–11, 2025; San Diego, CA, USAbooktitle: 54th International Conference on Parallel Processing (ICPP ’25), September 08–11, 2025, San Diego, CA, USAdoi: 10.1145/3754598.3754651isbn: 979-8-4007-2074-1/25/09ccs: Computing methodologies Concurrent algorithmsccs: Computing methodologies Massively parallel algorithmsccs: Computing methodologies Vector / streaming algorithms
1. Introduction
Distribution Statement A: Approved for public release. Distribution is unlimited.
Tensors provide a natural way to represent data with multiple dimensions. Tensor decomposition transforms tensors with higher dimensionalities to a reduced latent space that can be leveraged to learn salient features of the underlying data distribution. Different application domains, including machine learning (Ji et al., 2019), signal processing (Cichocki et al., 2015), and network analysis (Fernandes et al., 2021) have employed tensor decomposition to achieve superior performance compared to conventional approaches. Canonical Polyadic Decomposition (CPD) (Hong et al., 2020) has become the widely used tensor decomposition method, where the Matricized Tensor Times Khatri-Rao Product (MTTKRP) is the computational bottleneck (Kolda and Bader, 2009).
Since real-world tensors are generally sparse, tensor formats only keep nonzero tensor elements to reduce memory consumption (Wijeratne et al., 2021; Helal et al., 2021; Wijeratne et al., 2023b). There is a need to develop optimized sparse tensor data layouts that support the highly irregular data access patterns of MTTKRP while accessing input tensors and factor matrices (Wijeratne et al., 2021; Helal et al., 2021; Wijeratne et al., 2023b).
Real-world tensors often exhibit irregular shapes and distributions of nonzero tensor elements, which pose significant challenges when performing MTTKRP computations on multiple GPUs. These challenges arise from irregular memory access patterns, load imbalance among many GPU streaming multiprocessors, and the synchronization overhead among GPUs.
The size of real-world tensors is increasing rapidly with recent advances in big data applications (Cichocki, 2014). Such applications use tensors with billions of nonzero tensor elements (i.e., billion-scale tensors). Therefore, parallel algorithms that scale beyond a single GPU are required to perform MTTKRP (Bharadwaj et al., 2024; Chen et al., 2022a) on such large tensors. However, distributing the sparse tensor across multiple GPUs can lead to load imbalance, latency in data migration, and intermediate value communication among devices, which can result in additional overhead to execution time.
Recent works have proposed accelerating MTTKRP on CPU (Helal et al., 2021; Kurt et al., 2022; Wijeratne et al., 2023a; Laukemann et al., 2024), GPU (Nisa et al., 2019a, c; Wijeratne et al., 2024; Nguyen et al., 2022a; Lin et al., 2024), and ASIC (Srivastava et al., 2020). In CPU and ASIC, the external memory is large enough to maintain billion-scale tensors. Meanwhile, general-purpose GPUs encounter additional scalability challenges on a single device due to limited GPU global memory.
In prior work, multiple GPUs were used to accelerate MTTKRP on millions-scale sparse tensors (Chen et al., 2022b). Such implementations do not scale well on billion-scale sparse tensors because of (1) the significant idle time incurred by some GPUs due to workload imbalance across GPUs, (2) increased communication overhead due to the latency in sharing intermediate results across GPUs, (3) the synchronization overhead across GPUs, and (4) scheduling overhead of dynamic load-balancing schemes during execution.
Various efforts have been made to balance the workload of MTTKRP across streaming multiprocessors (SMs) of a single GPU (Nisa et al., 2019c; Wijeratne et al., 2024; Nguyen et al., 2022a; Li et al., 2018a). FLYCOO-GPU (Wijeratne et al., 2024) introduces a partitioning scheme that eliminates task dependencies between GPU SMs for each mode and minimizes workload imbalance among GPU SMs. Our work extends the FLYCOO-GPU (Wijeratne et al., 2024) partitioning scheme to a multi-GPU environment, addressing a new set of challenges: (1) eliminating data dependencies between tensor partitions to minimize GPU-GPU communication and synchronization overhead, (2) balancing the workload across GPUs to minimize the idle time for each GPU, and (3) avoiding host CPU computations when collecting and distributing input tensor partitions and factor matrices, since CPU computing power is significantly lower than GPUs.
Table 1. Summary of Related Work
Work Number of tensor Multi-GPU Load-balancing Support for Task independent
copies required support billion-scale tensors partitioning across GPUs
AMPED (ours) No. of modes
MM-CSF (Nisa et al., 2019c) No. of modes
FLYCOO-GPU (Wijeratne et al., 2024) 2
BLCO (Nguyen et al., 2022a) 1
HPSPTM (Chen et al., 2022b) No. of modes
ParTI-GPU (Li et al., 2018a) 1
*: Only support single GPU
The key contributions of this work are:
[…]
2. Background and Related Work
2.1. Introduction to Tensors
Table 2. Notations
Symbol Details
$\circ$ vector outer product
$\otimes$ Kronecker product
$\odot$ Khatri-Rao product
$\mathbf{A}$ matrix
$\mathbf{a}$ vector
$a$ scalar
$\mathcal{X}$ sparse tensor
$\mathcal{X}_{(d)}$ mode-$d$ matricization of $\mathcal{X}$
A tensor is a generalization of an array to multiple dimensions. In the simplest high-dimensional case, a tensor is a three-dimensional array, which can be visualized as a data cube. For a thorough review of tensors, refer to (Kolda and Bader, 2009). Table 2 summarizes the tensor notations.
2.1.1. Tensor mode
In Tensor Decomposition, the number of dimensions of an input tensor is commonly called the number of tensor modes. For example, a vector can be seen as a mode-1 tensor. A $N$-mode, real-valued tensor is denoted by $\mathcal{X}\in\mathbb{R}^{I_{0}\times\cdots\times I_{N-1}}$. This paper focuses on tensors of mode three or higher for tensor decomposition.
2.1.2. Indices of a nonzero tensor element
For a 3-mode tensor, $\mathcal{X}\in\mathbb{R}^{I_{0}\times I_{1}\times I_{2}}$, a nonzero tensor element is indicated as $x=\mathcal{X}(i_{0},i_{1},i_{2})$. Here, $i_{0}$, $i_{1}$, and $i_{2}$ are the positions or coordinates of $x$ in the tensor $\mathcal{X}$, which are commonly referred to as indices of the tensor element.
2.1.3. Tensor matricization
$\mathcal{X}_{(n)}$ denotes the mode-$n$ matricization or matrix unfolding (Favier and de Almeida, 2014) of $\mathcal{X}$. $\mathcal{X}_{(n)}$ is defined as the matrix $\mathcal{X}_{(n)}\in\mathbb{R}^{I_{n}\times(I_{0}\cdots I_{n-1}I_{n+1}\cdots I_{N-1})}$ where the parenthetical ordering indicates, the mode-$n$ column vectors are arranged by sweeping all the other mode indices through their ranges.
2.1.4. Canonical Poliyedic Tensor Decomposition (CPD)
CPD decomposes $\mathcal{X}$ into a sum of single-mode tensors (i.e., arrays), which best approximates $\mathcal{X}$. For example, given 3-mode tensor $\mathcal{X}\in\mathbb{R}^{I_{0}\times I_{1}\times I_{2}}$, our goal is to approximate the original tensor as $\mathcal{X}\approx\sum_{r=0}^{R-1}\mathbf{a}_{r}\circ\mathbf{b}_{r}\circ\mathbf{c}_{r}$, where $R$ is a positive integer and $\mathbf{a}_{r}\in\mathbb{R}^{I_{0}}$, $\mathbf{b}_{r}\in\mathbb{R}^{I_{1}}$, and $\mathbf{c}_{r}\in\mathbb{R}^{I_{2}}$.
For each of the three modes, the spMTTKRP operation can be expressed as
[…]
The alternating least squares (ALS) method is used to compute CPD. In a 3-mode tensor, CPD sequentially performs the computations in Equation 1, iteratively. This can be generalized to higher mode tensors. Note that the matricization of $\mathcal{X}$ is different for each factor matrix computation. In this paper, performing MTTKRP on all the matricizations of an input tensor is called computing MTTKRP along all the modes. The outputs $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$ are the factor matrices that approximate $\mathcal{X}$. $\mathbf{a}_{r}$, $\mathbf{b}_{r}$, and $\mathbf{c}_{r}$ refers to the $r^{\text{th}}$ column of $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$, respectively.
In this paper, we focus on MTTKRP on sparse tensors, which means that the tensor is sparse. Note that the factor matrices are dense.
2.2. Related Work
FLYCOO-GPU (Wijeratne et al., 2024) proposes a single GPU-based parallel algorithm to accelerate MTTKRP on sparse tensors while adopting the FLYCOO tensor format (Wijeratne et al., 2023b). FLYCOO-GPU (Wijeratne et al., 2023b) introduces dynamic tensor remapping on the GPU to reorder the tensor during execution time, allowing mode-specific optimizations. Dynamically remapping the input tensor limits the scalability of FLYCOO-GPU across GPUs as the inter-GPU communication becomes the bottleneck in multiple GPUs. Unlike (Wijeratne et al., 2024), our work supports multiple GPUs while avoiding inter-GPU task dependencies that lead to race conditions across GPUs.
BLCO (Nguyen et al., 2022a) proposes the blocked linearized coordinate format (BLCO) that enables out-of-memory computation where the input tensor is stored in host CPU external memory and streamed to a single GPU during the execution time of each mode computation. The streaming data-based computations of BLCO enable the partitions of the input tensor to be stored inside a large host CPU external memory, which is loaded into a single GPU over time and performs the MTTKRP partition by partition. Unlike BLCO, our approach distributes the computations across several GPUs while balancing the workload among all the GPUs.
HPSPTM (Chen et al., 2022b) proposes a framework to exploit multilevel parallelism and data reusability in heterogeneous HPC systems, including CPU + multi-GPU platforms. The data scheduling and partitioning scheme of HPSPTM dynamically transfers tensor partitions, intermediate results, and rows of factor matrices between GPUs during each factor matrix computation. In contrast, our work introduces a static load balancing scheme to minimize the data transfers across GPUs during each factor matrix computation. HPSPTM focuses on million-scale tensors, while our work focuses on billion-scale tensors.
MM-CSF (Nisa et al., 2019a) and HiCOO-GPU (Li et al., 2018a) propose novel tensor formats to distribute workload on a single GPU. In contrast, we propose a parallel algorithm to perform MTTKRP using multiple GPUs, which requires optimizations to reduce communication and synchronization overheads across GPUs.
Table 1 compares the characteristics of the related work with our work. Unlike related work, our work supports billion-scale sparse tensors. Our work also uses a task-independent partitioning scheme that balances the workload among the GPUs and their streaming multiprocessors.
3. Tensor Partitioning Scheme
In this work, we extend the partitioning scheme proposed in FLYCOO-GPU (Wijeratne et al., 2024) to distribute the nonzero tensor elements among multiple GPUs. Specifically, we adopt the FLYCOO-GPU partitioning scheme to distribute the tensor across GPUs, leveraging its task independence property, which allows each tensor partition to operate independently of the other partitions. For completeness, we describe the adopted partitioning scheme as inter-device partitioning.
Unlike FLYCOO-GPU (Wijeratne et al., 2024), our approach does not use dynamic tensor remapping. Consequently, we avoid embedding shard IDs within each nonzero tensor element. Instead, we maintain multiple copies of the input tensor in CPU external memory. For a detailed explanation of dynamic tensor remapping and the use of shard IDs, please refer to (Wijeratne et al., 2024).
Furthermore, we use the same notation introduced by FLYCOO-GPU (Wijeratne et al., 2024): When performing MTTKRP for the mode $d$ of an input tensor, we denote the mode $d$ as the output mode and its corresponding factor matrix as the output factor matrix. The rest of the tensor modes are called input modes, and the corresponding factor matrices are called input factor matrices.
3.0.1. Elementwise computation (EC)
Figure 1 summarizes the elementwise computation of a nonzero tensor element in mode 2 of a 3-mode tensor.
In Figure 1, the elementwise computation is carried out on a nonzero tensor element, denoted as $\mathcal{X}_{(2)}(i_{0},i_{1},i_{2})$. In sparse tensors, $\mathcal{X}_{(2)}(i_{0},i_{1},i_{2})$ is typically represented in formats such as COOrdinate (COO) format. These formats store the indices ($i_{0}$, $i_{1}$, and $i_{2}$) along with the element value (i.e., $val(\mathcal{X}_{(2)}(i_{0},i_{1},i_{2}))$).
To perform the computation, $\mathcal{X}_{(2)}(i_{0},i_{1},i_{2})$ is first loaded onto the processing units (i.e., streaming multiprocessors for GPU) from the external memory (step 1). The compute device retrieves the rows $\mathbf{{A}}(i_{0},:)$, $\mathbf{{B}}(i_{1},:)$, and $\mathbf{{C}}(i_{2},:)$ from the factor matrices using the index values extracted from $\mathcal{X}_{(2)}(i_{0},i_{1},i_{2})$ (step 2, step 3, and step 4). Then, the compute device performs the following computation:
$\mathbf{{C}}(i_{2},r)=\mathbf{{C}}(i_{2},r)+val(\mathcal{X}_{(2)}(i_{0},i_{1},i_{2}))\cdot\mathbf{{A}}(i_{0},r)\cdot\mathbf{{B}}(i_{1},r)$
Here, $r$ refers to the column index of a factor matrix row ($r<R$). The operation involves performing a Hadamard product between row $\mathbf{{A}}(i_{0},:)$ and row $\mathbf{{B}}(i_{1},:)$, and then multiplying each element of the resulting product by $val(\mathcal{X}_{(2)}(i_{0},i_{1},i_{2}))$. Finally, the updated value is stored in the external memory (step 5).
Figure 1. Elementwise computation (Wijeratne et al., 2024)
3.1. Tensor Partitioning
We introduce a static partitioning scheme to partition an input tensor along each mode of the input tensor. Similarly to related works (Nisa et al., 2019c, a), we use multiple copies of the input tensor, where each tensor copy is partitioned targeting an output mode computation. Note that all tensor copies are stored in the large external memory of the host CPU.
Our partitioning scheme in each mode distributes the nonzero tensor elements across (1) multiple GPUs using tensor sharding and (2) streaming multiprocessors (SMs) within each GPU using inter-shard partitioning. Note that a tensor shard is executed by a single GPU, and an inter-shard partition is executed by a single GPU SM.
3.1.1. Tensor Sharding
The proposed sharding scheme organizes the nonzero tensor elements of the input tensor into tensor shards (TS) based on the output mode index along each mode. For a given output mode $d$, the EC introduced in Section 3.0.1 is performed on each nonzero tensor element to update the corresponding rows of the output factor matrix. All nonzero tensor elements sharing the same output factor matrix index contribute to updating the same row. GPUs should maintain coherence among the updates that use nonzero tensor elements with the same output mode indices to avoid race conditions. To prevent race conditions between GPUs, all nonzero tensor elements sharing the same output mode index are assigned to the same TS, eliminating task dependencies across GPUs. It avoids the need to maintain coherency across GPUs during execution, thereby significantly reducing synchronization overhead.
3.1.2. Inter-Shard Partitioning
The nonzero tensor elements of each tensor shard (TS) are equally distributed among the SMs of each GPU by partitioning each TS into equal-sized inter-shard partitions (ISP). With ISP partitions, all the SMs of a GPU are assigned the same workload during execution. We use atomic operations to avoid race conditions between SMs of the same GPU.
3.2. Tensor Format Definition
Consider a multi-GPU setup with $m$ GPUs, each containing $g$ GPU SMs. For each output mode $d$, we divide the output mode indices $I_{d}$ into sets of equal-sized partitions $I_{d,0},I_{d,1},\ldots,I_{d,k_{d}-1}$, where $k_{d}=\frac{|I_{d}|}{m}$. Here, $|I_{d}|$ denotes the size of $I_{d}$. Each index partition $I_{d,j}$ ($j=0,1,\ldots,(k_{d}-1)$) is a subset of the output mode indices $I_{d}$. Next, all nonzero tensor elements that incident on the indices in $I_{d,j}$ are collected into a tensor shard, denoted by $\textit{TS}_{d,j}$.
To distribute nonzero tensor elements among GPU SMs, we further divide each tensor shard into equal-sized sets called inter-shard partitions (ISP). Each tensor shard $\textit{TS}_{d,j}$ is divided into $t_{d,j}=\left\lceil|\textit{TS}_{d,j}|/g\right\rceil$ inter-shard partitions. We denote the $q$-th inter-shard partition in $\textit{TS}_{d,j}$ as $\textit{ISP}_{d,j,q}$. For a tensor with $|\mathcal{T}|$ nonzero tensor elements, the total number of inter-shard partitions in mode $d$ is $\tau_{d}=\sum_{h=0}^{k_{d}-1}t_{d,h}\approx\frac{|\mathcal{T}|}{(m\times g)}$.
4. Parallel Algorithm
4.1. CUDA Programming Model
In the CUDA programming model (Owens et al., 2008; Hijma et al., 2023), a multi-threaded program is partitioned into blocks of threads (i.e., threadblocks) where each threadblock operates independently. The threadblocks are organized into a GPU Grid (Ansorge, 2022; Zeller, 2011). A multi-GPU implementation executes a kernel as Grids of threadblocks where each Grid is executed on a separate GPU. Each threadblock is executed by one GPU streaming multiprocessor (SM).
4.2. Tensor Partition and Threadblock Mapping
Figure 2. Tensor partition mapping of mode d
Following the partition scheme proposed in Section 3.1, the tensor shards (TS) and the inter-shard partitions (ISP) are assigned to GPU Grids and threadblocks, respectively (Ansorge, 2022), as illustrated in Figure 2. Figure 2 uses the same notation introduced in Section 3.1. Once a GPU finishes executing all the computations in a Grid, a new Grid is loaded onto the GPU for execution. Similarly, when a GPU SM finishes executing all the computations in a threadblock, a new threadblock from the same Grid is assigned to the SM.
4.3. Target Platform
Figure 3. Target platform
We consider a single node multi-GPU platform where the GPUs are connected to a single CPU, as illustrated in Figure 3. All the GPUs are connected using GPU to GPU interconnection (i.e., GPUDirect P2P (NVIDIA Corporation, 2024a)). Using the Cuda programming model, our parallel algorithm directly transfers data from one GPU global memory to another GPU global memory when required (see Section 4).
In Figure 3, we show GPU to GPU interconnection and host CPU to GPU interconnection separately for clarity. The host CPU to GPU interconnection and GPU to GPU interconnection can share the same physical hardware connections such as PCIe.
4.4. Data Distribution Among GPUs
All the tensor shards (for all output modes) are stored in the host CPU external memory. When a tensor shard ($TS$) is ready for execution, it is transferred to the global memory of the corresponding GPU.
The factor matrices of each billion-scale tensor are significantly smaller (i.e., a few megabytes) than the GPU global memory. Each GPU maintains a local copy of the factor matrices in its global memory. Once the MTTKRP computation for an output mode is completed, the updated rows of the output factor matrix are exchanged between GPUs to prepare for the computation of the next output mode.
4.5. Overall Algorithm
In this Section, we present the parallel algorithm for a single iteration of tensor decomposition. In tensor decomposition, the proposed algorithm is iteratively performed to generate the factor matrices that best approximate the original input tensor.
4.6. Mode-by-mode MTTKRP
Algorithm 1 shows the parallel algorithm for performing MTTKRP. Algorithm 1 takes (1) all the input tensor copies T and (2) factor matrices denoted as $\textbf{Y}=\{Y_{0},Y_{1},...,Y_{N-1}\}$. As shown in Algorithm 1, the MTTKRP is performed mode by mode (Algorithm 1: line 6). In each mode, A GPU grid (Algorithm 1: line 7) operates on a tensor shard mapped onto a GPU. At the end of all the computations of one mode, the GPUs are globally synchronized, and the generated output factor matrix rows are exchanged across GPUs before the computations of the next mode to maintain the correctness of the program (Algorithm 1: lines 8 - 11).
1 Input: Input tensor copies, $\textbf{T}=\{T_{0},T_{1}\cdots T_{N-1}\}$
2 Randomly initialized factor matrices, $\textbf{Y}=\{Y_{0},Y_{1},...,Y_{N-1}\}$
3 Output: Updated factor matrices $\hat{\textbf{Y}}=\{\hat{Y}_{0},\hat{Y}_{1},...,\hat{Y}_{N-1}\}$
4 for each mode $d=0,\ldots,N-1$ do
5 $T_{in}\leftarrow T_{d}$
6 // Execute Grids using multiple GPUs
7 for each tensor shard, $TS_{d,z}$ in $T_{in}$ parallel do
8    $\hat{Y}_{d}$ $\leftarrow$ GPU Grid($TS_{d,z}$, Y)
9
10 __Inter-GPU Barrier__
11 // Exchange generated output factor matrix partition across GPU
12 $Y_{d}$ $\leftarrow$ All Gather($\hat{Y}_{d}$)
13 __Inter-GPU Barrier__
14
Algorithm 1 Overall Proposed Algorithm
After processing TSTS, the updated factor matrix rows are shared across the GPUs (Algorithm 1: line 10) using all-gather communication primitive (Kobus et al., 2019) before executing MTTKRP for the next output mode.
4.7. Mapping Parallel Algorithm to GPU threadblocks
Figure 4. Overview of threadblock
1 GPU Grid($TS_{d,z}$, Y, $T_{out}$):
2 Input: Input tensor shard, $TS_{d,z}$
3 Factor matrices $\textbf{Y}=\{Y_{0},Y_{1},...,Y_{N-1}\}$
4
5Output: Updated factor matrix of mode $d$, $\hat{Y}_{d}$
6
7// Execute threadblocks in parallel across GPU SMs
8 for each inter-shard partition $ISP_{d,z,b}$ in $TS_{d,z}$, $b=0,1,\ldots,t_{d,z-1}$ parallel do
9 for each column, $t$ in threadblock parallel do
10    if $nnz+t<|ISP_{d,z}|$ then
11       Load($x_{i}$ at $(nnz+t)$)
12       $value\leftarrow val_{i}$
13       $p_{i}=(c_{0},\ldots,c_{N-1})$
14       // Elementwise Computation
15       for input mode $w\in\{0,\ldots,N-1\}\setminus\{d\}$ do
16
17          $vec\leftarrow$ Load(row $c_{w}$ from $w^{\text{th}}$ factor matrix)
18          // Row 0 to R1R-1 of the threadblock perform independent computations
19          for each rank $r$ in $R$ parallel do
20             $\ell(r)\leftarrow\ell(r)\times vec(r)$
21
22
23
24    for each rank $r$ in $R$ parallel do
25       $\hat{Y}_{d}(c_{d},r)\leftarrow\text{Atomic}(\hat{Y}_{d}(c_{d},r)+\ell(r))$
26
27 // P is the number of columns in a threadblock
28 $nnz\leftarrow nnz+P$
29
30
Algorithm 2 Parallel Algorithm Executed on a Shard
The basic computing unit of a GPU is a thread. According to the GPU programming model (Ansorge, 2022; Zeller, 2011), a multi-threaded program is partitioned into blocks of threads (i.e., threadblocks) that operate independently.
In our proposed algorithm, a threadblock has a dimension of $R\times P$, where $R$ denotes the rank of factor matrices and $P$ indicates the number of nonzero tensor elements parallelly loaded to a threadblock (see Figure 4). Here, each column of the threadblock shares the same nonzero tensor element, and each column performs elementwise computation on a nonzero tensor element.
Algorithm 2 outlines the computations executed on a shard partition. In Algorithm 2, $ISP_{d,z,p}$ corresponds to $p^{\text{th}}$ inter-shard partition in $z^{\text{th}}$ tensor partition of mode $d$. When a GPU SM is idle, a threadblock and its corresponding inter-shard partition are assigned to the GPU SM for computation. Each column in the threadblock loads a single nonzero tensor element at a time and shares the nonzero tensor element across the threads in the same column. Each thread in a column extracts the information from the tensor element $x_{i}$ in COO format (Algorithm 2: lines 9-11). Subsequently, each threadblock performs elementwise computation (Algorithm 2: lines 6-17). To achieve threadwise parallelism, each thread in a column only executes the update operation on a single column of a row of the output factor matrix (Algorithm 2: lines 15 - 17). The rows of the input factor matrices are loaded from the GPU global memory (Algorithm 2: lines 13-14) depending on the indices of the current tensor element ($p_{i}$) executed in the GPU thread. Each GPU threadblock locally updates the output factor matrix (Algorithm 2: lines 15) while maintaining the coherence of each threadblock to ensure the correctness of the program. According to the proposed partitioning scheme, threadblocks require atomic operations for tensor elements that reside in the same tensor shard. Hence, we use atomic operations across the GPU threadblocks within the same GPU to maintain the correctness of the program (Algorithm 2: lines 18-19).
4.8. Host CPU-GPU Communication
The input tensor copies are initially stored in the host CPU memory. The tensor shards are transferred from the host CPU memory to each GPU global memory during the execution time of each output mode.
4.9. All Gather Communication (GPU to GPU)
All GPUs maintain a local copy of the factor matrices in GPU global memory (see Section 4.4). As mentioned in Section 4.4, the factor matrices of each billion-scale tensor are significantly smaller (i.e., a few megabytes) than the GPU global memory. Hence, keeping a local copy of the factor matrices does not consume significant additional memory on each GPU.
For a given output mode, once a GPU completes processing the tensor shard (TS) assigned to it, the GPU communicates the updated rows to all the GPUs in the platform. The distributed factor matrix will act as an input factor matrix in the following modes of MTTKRP computations.
We adopt the ring network communication model (Li et al., 2018b) to perform all-gather among the GPUs, as shown in Algorithm 3. We use GPU-to-GPU communication to send and receive partitions of the factor matrices across GPUs without involving the CPU host memory. The ring network model is suitable for bulk transfers among neighboring devices with limited bandwidth, which is ideal for communicating the output factor matrix among the GPUs.
1 All Gather ($\hat{Y}_{d,gpu\_id}$):
2 Input: Mode $d$ output factor matrix partition in the local GPU, $\hat{Y}_{d,gpu\_id}$
3 // Using ring network communication model
4 //number of GPUs = $M$
5 for each GPU $z=0,\ldots,(M-2)$ do
6 // Send and Receive commands are executed in parallel
7   Send_Copy = $\hat{Y}_{d,(gpu\_id+z)modM}$
8   Send(Send_Copy, (gpu_id+1)modM(gpu\_id+1)modM, size(Send_Copy))
9   Rec_Copy = $\hat{Y}_{d,(gpu\_id-z-1)modM}$
10   Receive(Rec_Copy, (gpu_id1)modM(gpu\_id-1)modM, size(Rec_Copy)))
11   Barrier
12
Algorithm 3 GPU to GPU peer-to-peer communication
5. Experimental Results
5.1. Experimental Setup
5.1.1. Platform
Figure 5. Total execution time (speedup of our work compared to each baseline is shown at the top of each bar)
We conducted our experiments on a single CPU node with multi-GPUs. The CPU node has 4 NVIDIA RTX 6000 Ada Generation GPUs and AMD EPYC 9654 host CPU. The GPUs are connected to the host CPU via PCIe. GPUs use GPUDirect Peer-to-Peer (P2P) (NVIDIA Corporation, 2024a) to communicate with each other. Note that NVIDIA RTX 6000 Ada GPUs do not support NVLink (NVIDIA Corporation, 2024b) for direct GPU communication. Hence, we use the Direct P2P communication in our experiments.
GPU Specification: NVIDIA RTX 6000 Ada Generation GPU features the Ada Lovelace GPU architecture with 142 Streaming Multiprocessors (SMs) and 18176 cores, sharing 48 GB of GDDR6 global memory. Note that the NVIDIA RTX 6000 Ada Generation GPU has more computing power and global memory than the NVIDIA A100 GPU.
Host CPU Specification: In our experiments, we use a 2-socket AMD EPYC 9654 CPU as the host CPU platform. Each AMD EPYC 9654 consists of 96 physical cores (192 threads) running at a frequency of 2.4 GHz, sharing 1.5 TB of CPU external memory.
Each GPU is connected to the host CPU through a PCIe interface with 64 GB/s data bandwidth.
Table 3. Characteristics of the sparse tensors
Tensor Shape Number of
Tensor Elements
Amazon (Smith et al., 2017) $4.8M\times 1.8M\times 1.8M$ $1.7B$
Patents (Smith et al., 2017) $46\times 239.2K\times 239.2K$ $3.6B$
Reddit-2015 (Smith et al., 2017) $8.2M\times 177K\times 8.1M$ $4.7B$
Twitch (Rappaz et al., 2021; McAuley, 2021) $15.5M\times 6.2M\times 783.9K\times 6.1K\times 6.1K$ $0.5B$
5.1.2. Implementation
The source code is developed using CUDA C++ (Zeller, 2011) and compiled using CUDA version 12.2 (Fatica, 2008).
5.1.3. Datasets
We used all publicly available billion-scale tensors from the Formidable Repository of Open Sparse Tensors and Tools (FROSTT) dataset (Smith et al., 2017) and Recommender Systems and Personalization Datasets (Rappaz et al., 2021; McAuley, 2021). Table 3 summarizes the characteristics of the tensors.
5.1.4. Baselines
We evaluate the performance of our work by comparing it with the state-of-the-art GPU implementations, BLCO (Nguyen et al., 2022a) (out-of-memory computation enabled), MM-CSF (Nisa et al., 2019c), HiCOO-GPU (Li et al., 2018a), and FLYCOO-GPU (Wijeratne et al., 2024). The out-of-memory computation of BLCO stores the partitions of the input tensor inside the host CPU external memory, which is loaded into a single GPU over time and performs the MTTKRP partition by partition. To achieve optimal results with HiCOO-GPU, we use the recommended configurations provided in the source code (Li et al., 2019). For our experiments, we utilize the open-source BLCO repository (Nguyen et al., 2022b), ParTI repository (Li et al., 2019), and MM-CSF (Nisa et al., 2019b) repository.
5.1.5. Default Configuration
In our experiments, we used 4 GPUs as the default number of GPUs in the system. We set $\theta=32$ and $R=32$ as the configuration of the threadblocks described in Section 4.5.
5.1.6. Performance Metric - Total Execution Time
Similar to the literature (Nisa et al., 2019c; Wijeratne et al., 2024; Nguyen et al., 2022a; Li et al., 2018a), we measure the performance using the execution time to compute MTTKRP across all modes of an input tensor in a single iteration of tensor decomposition.
5.2. Overall Performance
Figure 5 shows the total execution time of our work on 4 NVIDIA RTX 6000 Ada GPUs. The speedup achieved by our work compared with each baseline in each input tensor is displayed at the top of the respective bar. The runtime error indicates that the host CPU operating system terminated the baseline during execution due to insufficient memory on the target hardware to store the input tensor, factor matrices and intermediate values.
[…]
When out-of-memory computation is enabled, BLCO (Nguyen et al., 2022a) stores the input tensor in the host CPU memory and loads the tensor partition by partition from the external memory of the host CPU to the GPU global memory. In BLCO, using a single GPU introduces additional memory traffic between the host CPU and the GPU. Meanwhile, our work has more effective bandwidth between the host CPU and the GPU since multiple GPUs can concurrently communicate with the host CPU. Our work shows a geometric speedup of $5.1\times$ using 4 GPUs compared to BLCO.
MM-CSF (Nisa et al., 2019c) performs MTTKRP only on the Amazon dataset. For Patents and Reddit, the GPU ran out of memory during the execution time. ParTI-GPU (Li et al., 2018a) can perform MTTKRP on Amazon and Patents. Also, MM-CSF (Nisa et al., 2019c) and ParTI-GPU (Li et al., 2018a) do not support Twitch, which has 5 modes.
[…]
5.3. Impact of Partitioning Scheme
Figure 6. Impact of proposed partitioning scheme
Section 3 describes the partitioning scheme used in this work to distribute the tensor shards among GPUs.
[…]
5.4. Execution Time Breakdown
For large tensors such as Patents and Reddit, moving tensor shards from CPU host memory to GPU global memory is the major contributor to the total communication time. Tensors with a large number of indices (e.g., Amazon and Twitch) require frequent GPU-GPU communication to update factor matrices at the end of each output mode computation, which significantly contributes to the total communication time. In particular, Reddit exhibits a significant communication overhead ($32\%$) during the execution due to (1) the large number of nonzero tensor elements in the tensor and (2) the modes with a large number of indices.
Figure 7. Execution time breakdown of the input tensors
5.5. Workload Distribution among GPUs
Figure 8. Computation time overhead among GPUs
As discussed in Section 4, each GPU updates a set of rows of the output factor matrix. In this section, the computation time of a GPU is defined as the total time to perform elementwise computation (EC) on all nonzero tensor elements assigned to a GPU across all modes. The computation time overhead is the difference between the maximum and minimum computation times among GPUs on the target platform.
Figure 8 shows the computation time overhead among GPUs (as a percentage) on our target 4 GPU platform. To determine the computation time of each GPU, we execute each GPU grid (see Section 4.2) separately and measure the computation time in each output mode. The computation time overhead for all billion-scale tensors is less than $1\%$ (as a percentage of the total time required to perform all EC using all 4 GPUs).
However, Twitch has the most computation time overhead due to some indices of the tensor corresponding to popular streamers and games in the Twitch platform, which leads to a disproportionately large number of nonzero tensor elements assigned to those indices. It results in a workload imbalance between GPUs, leading to a larger overhead.
5.6. Scalability
[…]
In the single GPU implementation, tensor partitions are stored in the host CPU external memory and loaded onto the GPU one iDp at a time in each mode.
Figure 9. Scalability of the proposed Algorithm
5.7. Preprocessing Time
The preprocessing of an input tensor involves generating tensor partitions following Section 3. Note that our work does not focus on the acceleration of preprocessing time. We have included a comparison of preprocessing times in Figure 10 for completeness. For comparison, we used the preprocessing time of BLCO. The host CPU described in Section 5.1.1 is used for preprocessing.
Figure 10. Preprocessing time
6. Conclusion and Future Work
[…]
In our future work, we will adapt the proposed parallel algorithm to heterogeneous computing platforms with different devices, such as multiple CPUs, GPUs, and Field Programmable Gate Arrays (FPGAs), to show the adaptability of the proposed parallel algorithm.
Acknowledgements.
This work is supported by the National Science Foundation (NSF) under grant OAC-2209563, CSSI-2311870, and in part by the DEVCOM Army Research Lab under grant W911NF2220159.